# Why Do We Laugh? Annotation and Taxonomy Generation for Laughable Contexts in Spontaneous Text Conversation

### v1

@@ Abstract
Laughter serves as a multifaceted communicative signal in human interaction, yet its identification within dialogue presents a significant challenge for conversational AI systems. This study addresses this challenge by annotating laughable contexts in Japanese spontaneous text conversation data and developing a taxonomy to classify the underlying reasons for such contexts. Initially, multiple annotators manually labeled laughable contexts using a binary decision (laughable or non-laughable). Subsequently, an LLM was used to generate explanations for the binary annotations of laughable contexts, which were then categorized into a taxonomy comprising ten categories, including “Empathy and Affinity” and “Humor and Surprise,” highlighting the diverse range of laughter-inducing scenarios. The study also evaluated GPT-4’s performance in recognizing the majority labels of laughable contexts, achieving an F1 score of 43.14%. These findings contribute to the advancement of conversational AI by establishing a foundation for more nuanced recognition and generation of laughter, ultimately fostering more natural and engaging human-AI interactions.
@@ 1 Introduction
In human dialogue, laughter serves as a communicative signal conveying humor, empathy, surprise, or social bonding Norrick (1993); Glenn (2003); Attardo (2009). However, its mechanisms are complex and multifaceted, and understanding them remains a long-term challenge for dialogue systems aiming to achieve human-like interaction Tian et al. (2016); Türker et al. (2017); Inoue et al. (2022); Ludusan and Wagner (2023); Perkins Booker et al. (2024). Furthermore, traditional approaches to modeling laughter and humor have often been limited to scenarios involving explicit auditory or visual stimuli, with few addressing the subtle contextual nuances present in spontaneous dialogue Bertero and Fung (2016); Choube and Soleymani (2020); Jentzsch and Kersting (2023); Ko et al. (2023); Hessel et al. (2023). Therefore, elucidating the underlying reasons for laughter in spontaneous dialogue data can contribute to making large language model (LLM)-based dialogue more natural and empathetic. However, annotating the reasons for laughter in any formalized manner has been prohibitively time- and labor-intensive, leaving the field largely reliant on qualitative approaches through conversational analysis.
@@
In this study, we address the question of “why do we laugh?” from an informatics perspective by proposing a semi-automated approach to constructing taxonomy labels for the reasons of laughter. First, to identify target segments, multiple annotators were asked to perform a simple binary classification on each utterance in a dialogue data, determining whether it was “laughable” or not, as shown in Table 1. Subsequently, for contexts labeled as “laughable” based on the majority voting, we used an LLM (GPT-4o) to generate the reasoning sentence behind this judgment and further classified these generated reasons into distinct categories (taxonomy labels). This semi-automated taxonomy generation approach is generalizable and can be particularly effective in scenarios where manual annotation is limited to simpler labels, such as emotion labeling.
@@ 3 Generating Taxonomy of Laughable Reason
Table 4: Laughable context recognition performance of GPT-4o on each taxonomy lable
@@
Next, we attempted to summarize the generated reasoning text for laughable contexts by following a taxonomy generation approach using LLMs Wan et al. (2024). First, we randomly divided the generated reasons into smaller portions, each representing 5% of the data. Starting with the initial one portion, we used GPT-4o again to generate taxonomy labels and explanations, manually validating them as needed. We then iteratively updated the taxonomy by having the LLM re-generate the taxonomy information based on the previous taxonomy information and the portion data, until all data was processed. This resulted in the generation of ten taxonomy labels, summarized in Table 3, including categories like (1) Empathy and Affinity and (2) Humor and Surprise.
After generating the taxonomy labels, we again used the LLM to assign labels to each reason sample, allowing for multiple labels per sample. The results of this labeling are summarized on the right side of Table 3. While dominant categories like (1) Empathy and Affinity were observed, a substantial number of samples were also labeled under other categories such as (4) Self-Disclosure and Friendliness and (5) Cultural Background and Shared Understanding. This varied distribution supports the validity of the generated taxonomy labels. A correlation matrix for the taxonomy labels is reported in Appendix A.
Finally, we investigated related conversational analysis studies, as listed on the right side of Table 3. These findings further support the explanatory power of our generated taxonomy labels within the framework of conversational analysis research.

### v2

@@ Abstract
Laughter serves as a multifaceted communicative signal in human interaction, yet its identification within dialogue presents a significant challenge for conversational AI systems. This study addresses this challenge by annotating laughable contexts in Japanese spontaneous text conversation data and developing a taxonomy to classify the underlying reasons for such contexts. Initially, multiple annotators manually labeled laughable contexts using a binary decision (laughable or non-laughable). Subsequently, an LLM was used to generate explanations for the binary annotations of laughable contexts, which were then categorized into a taxonomy comprising ten categories, including “Empathy and Affinity” and “Humor and Surprise,” highlighting the diverse range of laughter-inducing scenarios. The study also evaluated GPT-4o’s performance in recognizing the majority labels of laughable contexts, achieving an F1 score of 43.14%. These findings contribute to the advancement of conversational AI by establishing a foundation for more nuanced recognition and generation of laughter, ultimately fostering more natural and engaging human-AI interactions.
@@
 [^0] This paper has been accepted for presentation at International Workshop on Spoken Dialogue Systems Technology 2025 (IWSDS 2025) and represents the author’s version of the work.
@@ 1 Introduction
In human dialogue, laughter serves as a communicative signal conveying humor, empathy, surprise, or social bonding Norrick (1993); Glenn (2003); Attardo (2009). However, its mechanisms are complex and multifaceted, and understanding them remains a long-term challenge for dialogue systems aiming to achieve human-like interaction Tian et al. (2016); Türker et al. (2017); Mazzocconi et al. (2020); Inoue et al. (2022); Ludusan and Wagner (2023); Perkins Booker et al. (2024). Furthermore, traditional approaches to modeling laughter and humor have often been limited to scenarios involving explicit auditory or visual stimuli, with few addressing the subtle contextual nuances present in spontaneous dialogue Bertero and Fung (2016); Choube and Soleymani (2020); Jentzsch and Kersting (2023); Ko et al. (2023); Hessel et al. (2023). Therefore, elucidating the underlying reasons for laughter in spontaneous dialogue data can contribute to making large language model (LLM)-based dialogue more natural and empathetic. However, annotating the reasons for laughter in any formalized manner has been prohibitively time- and labor-intensive, leaving the field largely reliant on qualitative approaches through conversational analysis.
@@
In this study, we address the question of “why do we laugh?” from an informatics perspective by proposing a semi-automated approach to constructing taxonomy labels for the reasons of laughter. First, to identify target segments, multiple annotators were asked to perform a simple binary classification on each utterance in dialogue data, determining whether it was “laughable” or not, as shown in Table 1. Subsequently, for contexts labeled as “laughable” based on the majority voting, we used an LLM (GPT-4o) to generate the reasoning sentence behind this judgment and further classified these generated reasons into distinct categories (taxonomy labels). This semi-automated taxonomy generation approach is generalizable and can be particularly effective in scenarios where manual annotation is limited to simpler labels, such as emotion labeling.
@@ 3 Generating Taxonomy of Laughable Reason
Table 4: Laughable context recognition performance of GPT-4o on each taxonomy label
@@
We then aimed to summarize the generated reasoning texts for laughable contexts by applying a taxonomy generation approach using LLMs Wan et al. (2024). First, we randomly divided the generated reason samples into smaller subsets, each comprising roughly 5% of the samples. Starting with the first subset, we gave the reason sentences GPT-4o to generate initial taxonomy labels and those explanations, which we manually validated when necessary. We then iteratively refined the taxonomy by having the LLM update it based on the previous taxonomy and the reason sentence data from each new subset, continuing this process until all data were processed. This resulted in ten taxonomy labels, summarized in Table 3, including categories such as (1) Empathy and Affinity and (2) Humor and Surprise.
After generating these taxonomy labels, we used the LLM to assign them to each reason sample, allowing for multiple labels per sample. The labeling results are shown on the right side of Table 3. While some categories, such as (1) Empathy and Affinity, were predominant, many samples were also assigned to other categories, including (4) Self-Disclosure and Friendliness and (5) Cultural Background and Shared Understanding. This broad distribution across categories reinforces the validity of the generated taxonomy. A correlation matrix showing relationships between the taxonomy labels is provided in Appendix A. Finally, we reviewed related studies in conversational analysis, as listed on the right side of Table 3. These studies further substantiate the explanatory power of our taxonomy within the context of conversational analysis research.